OpenStudy (anonymous):

Is my answer right??Solve each system graphically. Be sure to check your solution. If a system has an infinite number of solutions, use set−builder notation to write the solution set. If a system has no solution, state this. 2x − 3y = 6, 3y − 2x = −6

OpenStudy (anonymous):

Equation 1: 2x - 3y = 6 3y = 2x - 6 y = (2/3)x - 2 Equation 2: 3y - 2x = -6 3y = 2x - 6 y = (2/3)x - 2 { (2/3)x - 2 |

OpenStudy (dumbcow):

you are correct, notice they are both the same line, same slope same y-intercept therefore infinite solutions

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2x − 3y = 6, add both equations − 2x + 3y = −6 0 = 0 { er yes it does!] ...so there are an infinite number of solutions to this simultaneous equation

OpenStudy (anonymous):

Thank you for the medal Donnice x